# Yang-Mills extension of the Loop Quantum Gravity-corrected Maxwell equations

### v2

@@ III The Noether Procedure as our path towards the non-Abelian scenario
This yields four Maxwell-type equations in the non-Abelian version with LQG contributions :
$2A_{\gamma}[\nabla\cdot\vec{E}_{h}+gf_{hij}\vec{E}_{i}\cdot\vec{A}_{j}]+g\theta_{3}l_{P}^{2}f_{hij}[2(\nabla^{2}\vec{E}_{i})\cdot\vec{A}_{j}+\vec{E}_{i}\cdot(\nabla^{2}\vec{A}_{j})+\nabla^{2}(\vec{A}_{i}\cdot\vec{E}_{j})-$
 $\;2\nabla[(\nabla\vec{A}_{i})\cdot(\nabla\phi_{j})]]+g\theta_{8}l_{P}f_{hij}[\nabla\cdot(\vec{A}_{i}\times\vec{\dot{A}}_{j})+\vec{A}_{i}\cdot(\nabla\times\vec{E}_{j})]=0,$ (29)
$A_{\gamma}[\nabla\times\vec{B}_{h}-gf_{hij}\phi_{i}(\vec{\dot{A}}_{j})-gf_{hij}\phi_{i}\vec{E}_{j}+gf_{hij}(\vec{B}_{i}\times\vec{A}_{j})]+\theta_{3}l_{P}^{2}\{2\nabla^{2}(\nabla\times\vec{B}_{h})-gf_{hij}[(\nabla^{2}\vec{E}_{i})\phi_{j}+$
$\;\frac{1}{2}\nabla^{2}(\phi_{i}\vec{E}_{j})+[(\nabla^{2}\vec{B}_{i})\times\vec{A}_{j}]-\frac{1}{2}\phi_{i}\nabla^{2}(\vec{\dot{A}}_{j})+\frac{1}{2}\vec{E}_{i}(\nabla^{2}\phi_{j})-\frac{1}{2}(\vec{\dot{A}}_{i})\cdot(\nabla^{2}\phi_{j})-\frac{1}{2}\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})-$
$\;\frac{1}{2}[\vec{B}_{i}\times(\nabla^{2}\vec{A}_{j})]-\nabla^{2}[\phi_{i}(\vec{\dot{A}}_{j})]]\}-\theta_{8}l_{P}\{2\nabla^{2}\vec{B}_{h}-\frac{g}{2}f_{hij}[\partial_{t}(\vec{A}_{i}\times\vec{A}_{j})-[(\vec{\dot{E}}_{i})\times\vec{A}_{j}]-2\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-$
@@
$\;\nabla[\nabla\cdot(\vec{A}_{i}\times\vec{A}_{j})]+(\nabla\times\vec{E}_{i})\phi_{j}]\}+2\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}\nabla\times\left[2(\vec{B}^{2}\vec{B}_{h})+3gf_{hij}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}\right]=\frac{\partial\vec{E}_{h}}{\partial t},$ (30)
$A_{\gamma}[(\nabla\cdot\vec{B}_{h})+gf_{hij}\vec{A}_{i}\cdot\vec{B}_{j}]-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\nabla\cdot[(\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})-3\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})]+\frac{g}{2}\theta_{8}l_{P}f_{hij}\nabla\cdot[(\vec{\dot{A}}_{i})\cdot\phi_{j}]+$
@@
 $\;9g\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}f_{hij}\nabla\cdot[(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}]=0,$ (31)
$A_{\gamma}[(\nabla\times\vec{E}_{h})-gf_{hij}\phi_{i}\vec{B}_{j}+gf_{hij}(\vec{A}_{i}\times\vec{E}_{j})]+\theta_{3}l_{P}^{2}\nabla^{2}(\nabla\times\vec{E}_{h})-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}[\vec{B}_{i}\cdot(\nabla^{2}\phi_{j})-$
$\;(\nabla^{2}\vec{E}_{i}\times\vec{A}_{j})]-\theta_{8}l_{P}[\nabla^{2}\vec{E}_{h}+gf_{hij}\nabla\cdot(\vec{E}_{i}\cdot\vec{A}_{j})-\frac{g}{2}f_{hij}\nabla\times(\vec{A}_{i}\times(\vec{\dot{A}}_{j}))]=-\frac{\partial\vec{B}_{h}}{\partial t}.$ (32)
@@
When considering the four non-Abelian Maxwell-type equations for massless spin-1 vector bosons in the context of LQG effects, it is important to notice that when the generators of the group $SU(N)$ become trivial ($G_{h}=0$), the structure constants of the group are taken $f_{hij}=0$, we re-obtains the LQG-corrected Maxwell equations, (3), (4), (5), and (6). Each term in $\theta^{i}$ contributes independently to the structure constants, and the term $\theta_{4}$, coupled with the nonlinear magnetic field, modifies the new magnetic field equations. Previously, the Eq. (31), known as Gauss Magnetism, did not include magnetic monopoles. However, this is no longer true in the non-Abelian formulation, as terms $\theta_{3}$, $\theta_{8}$, and $\theta_{4}$ independently contribute to this equation. It is conceivable, based on this interpretation, that magnetic monopoles might exist as the Planck scale, possibly associated to these parameters.
@@ IV Conclusive Considerations
The four modified Maxwell equations in the context of LQG and non-Abelian present significant results, as the dimensionless parameters of the LQG contribute to all equations. In the Abelian case, this does not happen, it is still possible to notice that the nonlinear parameter of the magnetic field in the Hamiltonian/Lagrangian is present in both equations, in which the magnetic field is the core of the discussion. Furthermore, this non-linear term, together with the LQG parameters and the other parameters that we have already mentioned, is present in the magnetic monopole equation. So, it is possible that on the Planck scale, these parameters can together yield magnetic monopole solutions. It is important to remark that we have not worked in the covariant formalism. The idea is to express in a manifest way field strengths and the gauge potentials interact. The developments proposed in this paper may open up the way for future developments of LQG effects in unexplored areas of physics, such as the investigation of LQG effects in both the electroweak theory and QCD. It is important to notice that effects arising from quantum gravitation may be found in these theories at high-energy scales. In particular, considering the electroweak sector, in our next project, we propose to pursue an investigation on the anomalous gauge-boson couplings - more specifically the anomalous (neutral) photon/Z-boson 3- and 4-vertices - searched for in the ATLAS and CMS experiments of the LHC, having in program to get new constraints on the LQG parameters from accelerator physics.
@@ V Acknowledgements
26/202.437/2024. We are also grateful to P. A. Lima
@@
Maxwell equations in presence of a background magnetic field.

### v3

@@ III The Noether Procedure as our path towards the non-Abelian scenario
From Eq. (29) and by taking its curl, it is possible to derive the Faraday-Lenz equation. Meanwhile, Gauss’s law for magnetism is obtained from Eq. (30) by taking the divergence of the magnetic field. This yields four Maxwell-type equations in the non-Abelian version including the LQG contributions :
$A_{\gamma}[\nabla\cdot\vec{E}_{h}+gf_{hij}\vec{E}_{i}\cdot\vec{A}_{j}]+\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}[2(\nabla^{2}\vec{E}_{i})\cdot\vec{A}_{j}+\vec{E}_{i}\cdot(\nabla^{2}\vec{A}_{j})+\nabla^{2}(\vec{A}_{i}\cdot\vec{E}_{j})$
$-2\nabla[(\nabla\vec{A}_{i})\cdot(\nabla\phi_{j})]]+\frac{g}{2}\theta_{8}l_{P}f_{hij}[\nabla\cdot(\vec{A}_{i}\times\vec{\dot{A}}_{j})+\vec{A}_{i}\cdot(\nabla\times\vec{E}_{j})]=0,$ (31)
$A_{\gamma}[\nabla\times\vec{B}_{h}+gf_{hij}\phi_{i}\vec{E}_{j}+gf_{hij}(A_{i}\times B_{j})]+\theta_{3}l_{P}^{2}\{2\nabla^{2}(\nabla\times\vec{B}_{h})-gf_{hij}[(\nabla^{2}\vec{E}_{i})\phi_{j}+\frac{1}{2}\nabla^{2}(\phi_{i}\vec{E}_{j})$
$+[(\nabla^{2}\vec{B}_{i})\times\vec{A}_{j}]-\frac{1}{2}\phi_{i}\nabla^{2}(\vec{\dot{A}}_{j})+\frac{1}{2}\vec{E}_{i}(\nabla^{2}\phi_{j})-\frac{1}{2}(\vec{\dot{A}}_{i})\cdot(\nabla^{2}\phi_{j})-\frac{1}{2}\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})-\frac{1}{2}[\vec{B}_{i}\times(\nabla^{2}\vec{A}_{j})]$
$-\nabla^{2}[\phi_{i}(\vec{\dot{A}}_{j})]]\}-\theta_{8}l_{P}\{2\nabla^{2}\vec{B}_{h}-\frac{g}{2}f_{hij}[\partial_{t}(\vec{A}_{i}\times\vec{A}_{j})-[(\vec{\dot{E}}_{i})\times\vec{A}_{j}]-2\nabla^{2}(\vec{A}_{i}\times\vec{A}_{j})-\nabla[\nabla\cdot(\vec{A}_{i}\times\vec{A}_{j})]$
@@
+(×$+(\nabla\times\vec{E}_{i})\phi_{j}]\}+2\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}\nabla\times\left[2(\vec{B}^{2}\vec{B}_{h})+3gf_{hij}(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}\right]=\frac{\partial\vec{E}_{h}}{\partial t},$ (32)
$A_{\gamma}[(\nabla\cdot\vec{B}_{h})+gf_{hij}\vec{A}_{i}\cdot\vec{B}_{j}]-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}\nabla\cdot[\vec{A}_{i}\times(\nabla^{2}\vec{A}_{j})-3\nabla^{2}(\vec{B}_{i}\times\vec{A}_{j})]+\frac{g}{2}\theta_{8}l_{P}f_{hij}\nabla\cdot[(\vec{\dot{A}}_{i})\cdot\phi_{j}]$
@@
$+9g\theta_{4}\mathcal{L}^{2}\left(\frac{\mathcal{L}}{l_{P}}\right)^{2\Upsilon_{\gamma}}l_{P}^{2}f_{hij}\nabla\cdot[(\vec{A}_{i}\times\vec{A}_{j})\cdot B^{2}]=0,$ (33)
$A_{\gamma}[(\nabla\times\vec{E}_{h})-gf_{hij}\phi_{i}\vec{B}_{j}+gf_{hij}(\vec{A}_{i}\times\vec{E}_{j})]+\theta_{3}l_{P}^{2}\nabla^{2}(\nabla\times\vec{E}_{h})-\frac{g}{2}\theta_{3}l_{P}^{2}f_{hij}[\vec{B}_{i}\cdot(\nabla^{2}\phi_{j})$
$-(\nabla^{2}\vec{E}_{i}\times\vec{A}_{j})]-\theta_{8}l_{P}[\nabla^{2}\vec{E}_{h}+gf_{hij}\nabla\cdot(\vec{E}_{i}\cdot\vec{A}_{j})-\frac{g}{2}f_{hij}\nabla\times(\vec{A}_{i}\times(\vec{\dot{A}}_{j}))]=-\frac{\partial\vec{B}_{h}}{\partial t}.$ (34)
@@
When considering the four non-Abelian Maxwell-type equations for massless spin-1 vector bosons in the context of LQG effects, it is important to notice that when the generators of the group $SU(N)$ become trivial ($G_{h}=0$), the structure constants of the group are taken $f_{hij}=0$, we re-obtains the LQG-corrected Maxwell equations, (3), (II), (5), and (6). Each term in $\theta^{i}$ contributes independently to the structure constants, and the term $\theta_{4}$, coupled with the nonlinear magnetic field, modifies the new magnetic field equations. Previously, the Eq. (33), known as Gauss Magnetism, did not include magnetic monopoles. However, this is no longer true in the non-Abelian formulation, as terms $\theta_{3}$, $\theta_{8}$, and $\theta_{4}$ independently contribute to this equation. It is conceivable, based on this interpretation, that magnetic monopoles might exist at the Planck scale, possibly associated to these parameters.
@@ IV Conclusive Considerations
The LQG-modified Maxwell and Yang-Mills equations open up new phenomenological prospects, as the dimensionless parameters of the LQG contribute to these equations and introduce tiny effects into Standard-Model processes. In the Abelian case, we highlight the presence os a non-linear contribution quadratic in the magnetic field. Furthermore, this non-linear term, along with the LQG and other parameters is present in the magnetic monopole equation. So, it is possible that at the Planck scale, these parameters may yield magnetic monopole solutions.
It is important to remark that we have not worked in the covariant formalism. The idea is to express in a manifest way field strengths and the gauge potentials interact. The developments proposed in this paper may open up the way for future developments of LQG effects in unexplored areas of physics, such as the investigation of LQG effects in both the electroweak theory and QCD. It is important to notice that effects arising from quantum gravitation may be found in these theories at high-energy scales. We are already pursuing the investigation of the effect of LQG corrections in connection with the electroweak sector of the Standard Model. More specifically, to our sense, it would be relevant to study the influence of LQG corrections on the anomalous vertices coupling the neutral gauge bosons, namely the photon and the Z0-boson. The present ATLAS and CMS experimental data could be used to get a new set of constraints on the LQG parameters. In particular, considering the electroweak sector, as mentioned, after the phase transition induced by the Higgs dublet, the third component of the SU(2) gauge boson and the weak hypercharge (Abelian) field get mixed. By performing the orthogonal transformation parametrized by the Weinberg angle, the photon and the Z0-boson field emerge as the physical fields, corresponding, respectively, to the zero mass and massive eigenstates of the mass matrix. Therefore, by introducing LQG effects in both the SU(2) and the U(1) factors, there will appear anomalous 3- and 4-vertices coupling directly the photon and the Z0-gauge boson. These vertices are searched for in the ATLAS and CMS Collaborations. Their present results may be used to establish a new bridge connecting LQG to accelerator physics.
@@ V Acknowledgements
26/202.437/2024. The authors are also grateful to P. A. Lima
@@
Maxwell equations in presence of background magnetic fields and astrophysical processes involving photons and neutrinos.
Appendix A From Maxwell to Yang-Mills: a glance at the Noether current procedure
Let us build up the Yang-Mills theory as a self-interacting theory of massless (gauge) vector bosons, starting from the ordinary Maxwell equations, as follows:
$\nabla\cdot\vec{E}=0,$ (35)
$\nabla\cdot\vec{B}=0,$ (36)
×$\nabla\times\vec{E}=-\frac{\partial\vec{B}}{\partial t},$ (37)
×$\nabla\times\vec{B}=\frac{\partial\vec{E}}{\partial t},$ (38)
The fields are represented in the the same form as in Eqs. (14), (15) and (16). Let us start off from the free lagrangian of the electromagnetic theory:
$\mathcal{L}_{EM}=\frac{1}{2}(E^{2}-B^{2}).$ (39)
Using the variational principle, it is possible to find both the time and space components of the conserved currents. The space component is given by
$\vec{j}_{h}=-f_{hij}[(\vec{E_{i}}\cdot\phi_{j})-(\vec{B_{i}}\times\vec{A_{j}})],$ (40)
whereas the time component reads as
$j_{h}^{0}=-f_{hij}(\vec{E_{i}}\cdot\vec{A_{j}}).$ (41)
Notice that the currents transform under the adjoint representation of the gauge group. So, they have a representation index different from the indices carried by the fields, since the latter are sitting in a general N-dimensional representation. To couple the currents to the potential fields in a gauge-invariant way, the two indices have to match. So, the gauge potentials must necessarily be put in the adjoint representation. In so doing, we can immediately couple the currents to the fields to get a new Lagrangian. Important: after the current/vector field coupling, the Lagrangian is no longer free; now, there appear self-interactions.
$\mathcal{L}_{EM}^{1}=\frac{1}{2}(E^{2}-B^{2})-l\phi_{h}j_{h}^{0}-l\vec{A_{h}}\vec{j_{h}},$ (42)
where $l$ is the free coupling constant parameter. We still have the dependence on space-time derivatives of the vector fields. To remove them, the field transformations are Eqs. (23), (24). From the Eq. (42), it is possible calculate the new field equations and the new on-shell currents associated to the Lagrangian:
$\vec{j_{k}^{(1)}}=-f_{khl}[(\vec{E}_{h}\cdot\phi_{l})-(\vec{B}_{h}\times\vec{A}_{l})-2lf_{hij}\vec{A}_{i}\phi_{l}\phi_{j}],$ (43)
$j_{k}^{0(1)}=-f_{khl}(\vec{E}_{h}\cdot\vec{A}_{l}-2lf_{hij}\vec{A}_{i}\vec{A}_{l}\phi_{j}).$ (44)
These two currents above signal the final step of ther procedure; so, we can couple them through the Lagrangian
$\mathcal{L}_{EM}^{2}=\mathcal{L}_{EM}^{1}-l^{\prime}\phi_{k}j_{k}^{0(1)}-l^{\prime}\vec{A_{k}}\vec{j}_{k}^{(1)},$ (45)
Again, the $l^{\prime}$ is another free coupling constant parameter. From the Eq. (45), it is possible obtain the equation of the electric and magnetic field in the Yang-Mills formalism
$\vec{E}_{h}=-\nabla\cdot\phi_{h}-\frac{\partial\vec{A}_{h}}{\partial t}+gf_{hij}\vec{A}_{i}\phi_{j},$ (46)
$\vec{B}_{h}=(\nabla\times\vec{A}_{h})+\frac{1}{2}gf_{hij}(\vec{A}_{i}\times\vec{A}_{j}).$ (47)
It is possible to obtain the Gauss’ electricity and Ampère-Maxwell equations from Eq. (45). However another strategy consists of obtaining the four non-Abelian Maxwell equations from the general form of the field Eqs. (46) and (47). If we take the curl and div from both equations, we can obtain the new ”non-Abelian” Maxwell equations: actually, the Yang-Mills equations.
$\nabla\cdot\vec{E}_{h}+gf_{hij}\vec{A}_{i}\vec{E}_{j}=0,$ (48)
$\nabla\cdot\vec{B}_{h}+gf_{hij}\vec{A}_{i}\vec{B}_{j}=0,$ (49)
×$\nabla\times\vec{E}_{h}+gf_{hij}(\vec{A}_{i}\times\vec{E}_{j})=-\frac{\partial\vec{B}_{h}}{\partial t}+gf_{hij}\phi_{i}\vec{B}_{j},$ (50)
×$\nabla\times\vec{B}_{h}+gf_{hij}(\vec{A}_{i}\times\vec{B}_{j})=\frac{\partial\vec{E}_{h}}{\partial t}-gf_{hij}\phi_{i}\vec{E}_{j}.$ (51)